A 500g slab of cheese is completely eaten by two mice. If Minnie had eaten twice as much as she did, and Mickey had eaten half as much as he did, there would have been 100g left. If Minnie = x and Mickey = y, use the information to get the 2 equations relating to x and y.
Answer by venugopalramana(3286)

A 500g slab of cheese is completely eaten by two mice. If Minnie had eaten twice as much as she did, and Mickey had eaten half as much as he did, there would have been 100g left. If Minnie = x and Mickey = y, use the information to get the 2 equations relating to x and y.
TOTAL EATEN BY MINNIE AND MICKEY =
X+Y = 500.........................1
MINNIE ATE =X
IF SHE HAD EATEN TWICE IT WOULD BE =2X
MICKEY ATE =Y
IF HE HAD EATEN HALF OF IT ONLY IT WOULD BE = Y/2
TOTAL =
2X+(Y/2)=500-100=400
4X+Y=800...............................2
SOLVING
EQN.2-EQN.1
3X=300
X=100
Y=400